Perez Electronics is considering investing in manufacturing equipment expected to cost $310,000. The equipment has an estimated useful life of four years and a salvage value of $ 18,000. It is expected to produce incremental cash revenues of $155,000 per year. Perez has an effective income tax rate of 40 percent and a desired rate of return of 10 percent. (PV of $1 and PVA of $1) (Use appropriate factor(s) from the tables provided.)

a.Determine the net present value and the present value index of the investment, assuming that Perez uses straight-line depreciation for financial and income tax reporting.

b.Determine the net present value and the present value index of the investment, assuming that Perez uses double-declining-balance depreciation for financial and income tax reporting.

d.Determine the payback period and unadjusted rate of return (use average investment), assuming that Perez uses straight-line depreciation.

e.Determine the payback period and unadjusted rate of return (use average investment), assuming that Perez uses double-declining-balance depreciation. (Note: Use average annual cash flow when computing the payback period and average annual income when determining the unadjusted rate of return.)
Answer #1

Answer a)

Year Cash revenue Depreciation EBT Tax (40%) EAT Cash flows (EAT+ DEP.)
1 155,000 73000 82000 32800 49200 122200
2 155,000 73000 82000 32800 49200 122200
3 155,000 73000 82000 32800 49200 122200
4 155,000 73000 82000 32800 49200 122200
Straight line Deprication = 310000-18000/4 =73000
Year Cash flow Present value @ 10 Present value of cash flow
1 122200 0.91 111079.80
2 122200 0.83 100937.20
3 122200 0.75 91761.09
4 122200 0.68 83419.17
Total P.V. 387197.26
Cost of Project 310,000
NPV 77197.26

Answer b)  

1st Year   -     Straight-line Depreciation Rate = 1 ÷ 4 = 0.25 = 25%
Declining Balance Rate = 2 × 25% = 50%
Depreciation = 50% × $310,000 = $155,000

2nd Year   Declining Balance Rate = 50%
Book Value = Cost ? Accumulated Depreciation = $310,000 ? $155,000 = $155000
Depreciation = 50% × $155,000 = $77500

3 rd Year   Declining Balance Rate = 50%
Book Value = Cost ? Accumulated Depreciation = $$155,000 ? $77500= $77500
Depreciation = 50% × $77500= $38750

4th Year   Declining Balance Rate = 50%
Book Value = Cost ? Accumulated Depreciation = $ 77500? $ 38750= $38750

Depreciation = 50% × $38750= $19375

Year Cash revenue Depreciation EBT Tax (40%) EAT Cash flows (EAT+ DEP.)
1 155,000 155,000 0 32800 -32800 40200
2 155,000 77500 77500 32800 44700 117700
3 155,000 38750 116250 32800 83450 156450
4 155,000 19375 135625 32800 102825 175825
Year Cash flow Present value @ 10 Present value of cash flow
1 40200 0.91 36541.80
2 117700 0.83 97220.20
3 156450 0.75 117479.73
4 175825 0.68 120025.99
Total P.V. 371267.72
Cost of Project 310,000
NPV 61267.72

Answer d)

Year Cash revenue Depreciation EBT Tax (40%) EAT Cash flows (EAT+ DEP.)
1 155,000 73000 82000 32800 49200 122200
2 155,000 73000 82000 32800 49200 122200
3 155,000 73000 82000 32800 49200 122200
4 155,000 73000 82000 32800 49200 122200

Payback period for 2 years = 122200+122200 = 244400

Cash flow for the 3rd year = 122200

Payback period = 2 years +310000-244400/122200*12= 6.4 months

Total payback period 2 years and 6.4 mn.

Answer e)

Year Cash revenue Depreciation EBT Tax (40%) EAT Cash flows (EAT+ DEP.)
1 155,000 155,000 0 32800 -32800 40200
2 155,000 77500 77500 32800 44700 117700
3 155,000 38750 116250 32800 83450 156450
4 155,000 19375 135625 32800 102825 175825

Payback period for 2 years = 40200+117700= 157900

Cash flow for the 3rd year = 156450

Payback period = 2 years + 310000-157900/156450*12= 11.7 months

Total payback period is 2 years and 11.7 mn